i need help with 11.89 a and b
(d) What is the standard deviation of the number of cars sold in a day? Table 11.14 Cars sold in a day 0 3 4 Plx) 0.29 0.3 0.2 0.17 11.89 More Fruit Fly Lifetimes Refer to Table 11.13 in Exercise 11.87 that gives probabilities for fruit fly lifetimes. (a) If we know a fruit fly died before the end of its second month, what is the probability it died in its first month? (b) If a fruit fly makes it past its second month what is the probability it will live more than four months? 11.90 More Used Car Sales Refer to Table 11.14 in Exercise 11.88 that gives probabilities for the num ber of used cars a dealer sells in a day. What is the probability that the dealership sells no cars on three consecutive days? (Assume daily sales are independent.) 11.91 Ray Allens Free Throws As we see in
media%2F2a4%2F2a46cdfa-be26-46ad-8019-74
Answer #1

total probability should be 1 ,

so the proportion of dying in the second month is

P(X = 2) = 1−(0.30 + 0.20 + 0.15 + 0.10 + 0.05) = 1−0.80 = 0.20

a)

P(die in first month | given that died before end of 2nd month) = P(X=1 | X≤2) = P(X=1)/P(X≤2) = 0.30/(0.30+0.20) =0.6

b)

P(X>4 | X≥2) = P(X>4)/P(X≥2) =(0.10+0.05)/(0.20+0.20+0.15+0.10+0.05)=0.2143